# Predictive Uncertainty Quantification via Risk Decompositions for Strictly Proper Scoring Rules

### v2

@@ Appendix D Relations between the estimates
$\tilde{\text{R}}_{\text{Tot}}^{(2)}(x)=\tilde{\text{R}}_{\text{Bayes}}^{(2)}(x)+\tilde{\text{R}}_{\text{Exc}}^{(2,2)}(x)=\tilde{\text{R}}_{\text{Bayes}}^{(1)}(x)+\tilde{\text{R}}_{\text{Exc}}^{(1,2)}(x).$
Additionally, Bregman information can be received as follows:
$BI(x)=\tilde{\text{R}}_{\text{Exc}}^{(1,1)}(x)-\tilde{\text{R}}_{\text{Exc}}^{(2,1)}(x)=\tilde{\text{R}}_{\text{Exc}}^{(1,2)}(x)-\tilde{\text{R}}_{\text{Exc}}^{(2,2)}(x)=\tilde{\text{R}}_{\text{Bayes}}^{(2)}(x)-\tilde{\text{R}}_{\text{Bayes}}^{(1)}(x).$
Reverse Bregman Information:
$RBI(x)=\tilde{\text{R}}_{\text{Exc}}^{(2,1)}(x)-\tilde{\text{R}}_{\text{Exc}}^{(2,2)}(x)=\tilde{\text{R}}_{\text{Exc}}^{(1,1)}(x)-\tilde{\text{R}}_{\text{Exc}}^{(1,2)}(x)=\tilde{\text{R}}_{\text{Tot}}^{(1)}(x)-\tilde{\text{R}}_{\text{Tot}}^{(2)}(x).$
Appendix E Hand-crafted datasets
In this section, we describe the noisy versions of CIFAR10 and CIFAR100 datasets created for our experiments. We created three noisy datasets, each discussed below.
E.1 Noisy labels datasets
In the dataset, the images are the same as in the original dataset (covariates are not changed). However, some of the labels are randomly swapped. Hence, only labels were changed, while covariates were kept as in the original dataset. The motivation for the creation of this dataset is due to the fact that conventional image classification datasets essentially contain no aleatoric uncertainty [19]. To mitigate the limitation, which is critical for our evaluation, we introduce the label noise manually. By nature, this noise is aleatoric.
CIFAR10.
CIFAR100.
E.2 Blurred datasets
When predicting by a model trained to particular data, on some another dataset, where covariates are drastically different (“hard-OOD”), predictions are not meaningful anymore.
To introduce “soft-OOD” datasets, where labels remain the same but covariates are altered (yet the predicted vectors are still reasonable), we applied a Gaussian blur to the images in CIFAR10 and CIFAR100. Specifically, before the same transformations that we did for standard validation splits, we did Gaussian blur with kernel size of $(3,3)$ and sigma of $(0.1,2.0)$. These blurred datasets help model “soft-OOD” scenarios.
E.3 Missing class dataset
Essentially, Excess risk-based uncertainty measures evaluate disagreement between models in an ensemble. It is interesting to study this disagreement in the case when some of the classes are missing for some members within the ensemble. For this, we consider a special version of the CIFAR10 dataset, when we totally ignore some classes for some members of the ensemble. It results in an increase disagreement between ensemble members.
In our experiments, we had overall 20 ensemble members. For members 1 and 2 class 0 was missed. For members 3 and 4 class 1 was missed. And for members 5 and 6 class 2 was missed. All other members of the ensemble had all the classes.
Appendix F Training details
Training procedures for each dataset were similar. We used either ResNet18 or VGG19 architectures.
In Figure 4 we present performance summary statistics of the ensembles of different architectures. Specifically, we show accuracy, macro averaged precision, recall, and F1-score. In the left figure, the distribution for ResNet18, trained with different loss functions. In the right figure, VGG19.
Similarly to CIFAR10, in Figure 5 we present performance summary statistics of the ensembles of different architectures. In the left figure, the distribution for ResNet18, trained with different loss functions. In the right figure, VGG19.
Figure 4: Violin plots for different training loss functions and different metrics. The training dataset is CIFAR10. Left: ResNet18. Right: VGG-19.
Figure 5: Violin plots for different training loss functions and different metrics. The training dataset is CIFAR100. Left: ResNet18. Right: VGG-19.
Appendix G Additional experiments on matching / not-matching proper scoring rules
Figure 6: AUROC for out-of-distribution detection. Left: When scoring rules, used for loss function and risks computation, match. Right: When they are not matching. All proper scoring rules are considered.
Figure 7: AUROC for out-of-distribution detection. Left: When scoring rules, used for loss function and risks computation, match. Right: When they are not matching. Neglog proper scoring rule excluded from consideration.
In this section, we provide additional results on the matching and non-matching proper scoring rules used for loss function and risk (uncertainty measures) computations.
Figure 6 displays a histogram of AUROC scores for out-of-distribution detection, comparing the matching and non-matching cases. For the matching case, the results are meaningful, showing a significant margin from an AUROC score of 0.5, which corresponds to a constant predictor. In contrast, the non-matching case includes values less than 0.5, indicating a drawback in combining different loss functions and uncertainty measures.
From Figure 2 in the main text, we see that the most unstable scoring rule is Neglog. One might think this instability is due to its computational issues, and excluding it might equalize the results for matching and non-matching cases. To test this, we plot an additional figure (Figure 7) excluding the Neglog scoring rule.
However, even after excluding Neglog, we still observe noise in the results, particularly from combinations of the Spherical loss function with other scoring rules for risk computations. This demonstrates that the effect of matching and non-matching scoring rules persists beyond the instability of any single rule.
Appendix H Additional experiments on out-of-distribution detection
In this section, we provide an additional comparison of matching and not-matching results for out-of-distribution detection. For this, we add Table 4 and compare columns “Matching” and “Not-Matching” in this table element-wise. Column “Matching” corresponds to matching proper scoring rule used for loss function and uncertainty measures. Column “Not-Matching” corresponds to different ones. We see, that results in the left part of Table 4, except two values (in red), are systematically better, than in its right part. However, those two values are very close to each other.
Dataset Matching Not-Matching
Bayes Excess Total Bayes Excess Total
InD OOD
CIFAR10 Blurred CIFAR10(*) 84.38 87.80 85.90 83.91 86.59 86.00
Blurred CIFAR100 94.48 95.65 95.42 94.25 94.95 95.34
CIFAR100 91.05 90.08 90.87 90.96 88.21 90.77
SVHN 94.54 93.31 94.44 94.46 91.90 94.30
CIFAR100 Blurred CIFAR10 87.96 85.36 90.32 87.65 76.68 87.96
Blurred CIFAR100(*) 71.40 77.55 74.15 70.99 72.83 73.55
CIFAR10 79.35 72.70 79.27 78.87 64.19 76.64
SVHN 84.90 73.99 84.59 84.91 65.46 82.12
Table 4: AUROC for OOD detection. Left (Matching): For computation of OOD measures, we used loss and corresponding risks, generated by the matching function $G$. Right (Not-matching): Loss and corresponding risks generated by not-matching functions $G$. The best results (element-wise, between Matching and Not-matching) are in bold. In red are those where Different is better. By asterisk (*) we denote “soft-OOD”.
Appendix I Additional experiments misclassification detection
Similarly to the previous section with out-of-distribution detection, we compare columns “Matching” and “Not-Matching” in Table 5 element-wise. Column “Matching” corresponds to matching proper scoring rule used for loss function and uncertainty measures. Column “Not-Matching” corresponds to different ones. We see, that results in the left part of Table 5, are systematically better, than in its right part.
Dataset Matching Not-Matching
Bayes Excess Total Bayes Excess Total
CIFAR10 94.53 94.65 94.78 94.30 92.72 94.39
CIFAR100 86.47 82.90 86.83 85.57 71.62 83.13
Missed class CIFAR10 93.73 83.23 91.14 93.55 79.71 89.75
Noisy CIFAR10 81.00 74.35 80.97 79.49 70.17 78.52
Noisy CIFAR100 82.64 72.30 82.45 81.49 60.23 78.19
Table 5: AUROC for misclassification detection. Left (Matching): For computation of OOD measures, we used loss and corresponding risks, generated by the matching function $G$. Right (Not-Matching): Loss and corresponding risks generated by not-matching functions $G$. The best results (element-wise, between Matching and Not-Matching) are in bold.
Appendix J Other approximations
In this section, we discuss choices of Bayesian approximation strategy (Inner or Outer) separately for different problems. Outer in this notation means the first approach of the Bayesian approximation, when we perform Bayesian averaging of risk, and effectively apply an outer expectation. Inner corresponds to the second approach of Bayesian approximation, namely Bayesian model averaging when we first find the expectation and then plug it in the risk. Effectively we do inner expectation in this case.
J.1 Out-of-distribution detection
Here, we extend the results presented in the main part, and demonstrate, how the approximation strategy of risks influences the results of out-of-distribution detection. We present results in Table 6, where matching scoring rules were used for loss function and risks computations, and Table 7, where they are different.
We see, that results in Table 6 (matching scoring rules) are systematically better, than results in Table 7, when there was no matching.
Bayes risk.
We see, that there is no the only leader of specific approximation. Nevertheless, in most cases and on average, Inner approximations seem slightly better. Average values for matching case: Bayes(O) - $85.47\pm 8.12$, Bayes(I) - $86.54\pm 7.69$. Average values for not matching case: Bayes(O) - $85.09\pm 8.32$, Bayes(I) - $86.41\pm 7.72$. In both cases, the Inner approximation is slightly (on average) better, than the Outer.
Total risk.
Similarly to Bayes risk, we see that measures are very close to each other. Average values for matching case: Total(O) - $87.20\pm 7.17$, Total(I) - $86.54\pm 7.69$. Average values for not matching case: Total(O) - $85.26\pm 8.37$, Total(I) - $86.41\pm 7.72$.
Excess risk.
In these Tables we have three out of four different approximations of Excess risk: Bregman Information (BI), Reverse Bregman Information (RBI), and Expected Pairwise Bregman Divergence (EPBD). We do not consider Inner Inner approximation, as it is zero. As in the previous cases, there is no sole leader in these measures.
Dataset Metrics
Bayes(O) Bayes(I) Total(O) Total(I) BI RBI EPBD
InD OOD
CIFAR10 Blurred CIFAR10(*) 83.26 85.49 86.30 85.49 87.63 87.93 87.84
Blurred CIFAR100 93.75 95.22 95.62 95.22 95.60 95.68 95.66
CIFAR100 91.11 90.99 90.75 90.99 90.20 89.96 90.07
SVHN 94.55 94.54 94.34 94.54 93.39 93.23 93.31
CIFAR100 Blurred CIFAR10 86.74 89.18 91.45 89.18 85.83 84.48 85.76
Blurred CIFAR100(*) 70.18 72.62 75.68 72.62 76.76 78.03 77.87
CIFAR10 79.25 79.44 79.10 79.44 73.49 71.82 72.78
SVHN 84.95 84.85 84.33 84.85 75.07 72.80 74.11
Table 6: AUROC for out-of-distribution detection, when matching scoring rules were used for loss/risks. Here we extend different approximations. Notation: O - Outer, I - Inner. (R)BI - (Reverse) Bregman Information, EPBD - Expected Pairwise Bregman Divergence. Best results in bold, second-best underline. By asterisk (*) we denote “soft-OOD”.
Dataset Metrics
Bayes(O) Bayes(I) Total(O) Total(I) BI RBI EPBD
InD OOD
CIFAR10 Blurred CIFAR10(*) 82.56 85.26 86.74 85.26 86.57 86.46 86.73
Blurred CIFAR100 93.45 95.05 95.64 95.05 95.00 94.71 95.12
CIFAR100 90.92 90.99 90.55 90.99 88.34 88.04 88.25
SVHN 94.41 94.51 94.09 94.51 91.94 91.75 92.00
CIFAR100 Blurred CIFAR10 86.36 88.94 86.98 88.94 76.98 75.04 78.02
Blurred CIFAR100(*) 69.42 72.56 74.54 72.56 71.87 72.99 73.63
CIFAR10 78.71 79.03 74.26 79.03 64.67 63.30 64.61
SVHN 84.87 84.95 79.29 84.95 66.20 64.34 65.84
Table 7: AUROC for out-of-distribution detection, when not-matching scoring rules were used for loss/risks. Notation: O - Outer, I - Inner. (R)BI - (Reverse) Bregman Information, EPBD - Expected Pairwise Bregman Divergence. Best results in bold, second-best underline. By asterisk (*) we denote “soft-OOD”.
J.2 Misclassification detection
Dataset Metrics
Bayes(O) Bayes(I) Total(O) Total(I) BI RBI EPBD
InD
CIFAR10 94.30 94.76 94.79 94.76 94.68 94.60 94.65
CIFAR100 86.12 86.81 86.86 86.81 83.74 81.92 83.04
Missed class CIFAR10 94.47 92.99 89.30 92.99 86.34 80.86 82.49
Noisy CIFAR10 80.57 81.42 80.53 81.42 74.77 73.98 74.31
Noisy CIFAR100 82.66 82.62 82.28 82.62 73.28 71.25 72.37
Table 8: AUROC for misclassification detection, when matching scoring rules were used for loss/risks. Notation: O - Outer, I - Inner. (R)BI - (Reverse) Bregman Information, EPBD - Expected Pairwise Bregman Divergence. Best results in bold, second-best underline.
Dataset Metrics
Bayes(O) Bayes(I) Total(O) Total(I) BI RBI EPBD
InD
CIFAR10 94.11 94.50 94.29 94.50 92.53 92.59 93.04
CIFAR100 85.19 85.96 80.30 85.96 72.45 70.00 72.40
Missed class CIFAR10 94.28 92.83 86.67 92.83 81.24 77.62 80.26
Noisy CIFAR10 78.52 80.45 76.59 80.45 71.02 69.18 70.33
Noisy CIFAR100 81.44 81.54 74.85 81.54 61.37 58.88 60.45
Table 9: AUROC for misclassification detection, when not-matching scoring rules were used for loss/risks. Notation: O - Outer, I - Inner. (R)BI - (Reverse) Bregman Information, EPBD - Expected Pairwise Bregman Divergence. Best results in bold, second-best underline.
Similar to the previous section, here we demonstrate extended results of misclassification detection. We present results in Table 8, where matching scoring rules were used for loss function and risks computations, and Table 9, where they are different.
We see, that results in Table 8 (matching scoring rules) are systematically better than results in Table 9, when there was no matching.
Bayes risk.
As before, there is no sole leader. Average values for matching case: Bayes(O) - $87.62\pm 6.48$, Bayes(I) - $87.72\pm 5.99$. Average values for not matching case: Bayes(O) - $86.71\pm 7.23$, Bayes(I) - $87.06\pm 6.40$.
Total risk.
Similarly to Bayes risk, we see that measures are very close to each other. Average values for matching case: Total(O) - $86.75\pm 5.70$, Total(I) - $87.72\pm 5.99$. Average values for not matching case: Total(O) - $82.54\pm 7.97$, Total(I) - $87.06\pm 6.40$.
Excess risk.
In these Tables we have three out of four different approximations of Excess risk: Bregman Information (BI), Reverse Bregman Information (RBI), and Expected Pairwise Bregman Divergence (EPBD). We do not consider Inner Inner approximation, as it is zero. As in the previous cases, there is no sole leader in these measures.

### v3

@@ I.2 ImageNet datasets
$\beta=\beta_{prior}+n-x,$
where by subscript “prior” we highlight the parameters of the prior distribution, $x$ is the number of successes in observed training dataset, and $n$ is the total number of observations.
Therefore we see, that given larger training dataset, the estimates of epistemic uncertainty will shrink. This is the property we typically want to have from the epistemic uncertainty estimates – the bigger the dataset, the less uncertainty.
However, in case of the prior misspecification (will be discussed below), it leads to certain problems. If the prior parameters $\alpha_{prior}$ and $\beta_{prior}$ are already large (indicating a highly concentrated prior), the posterior will also be highly concentrated, regardless of the data. This can lead to an underestimation of epistemic uncertainty because the model appears more certain than it should be.
M.1 Various distribution shapes
In this section, we explore different choices of the parameters and demonstrate, how the resulting posterior distribution will look like, given different prior and different sizes of training datasets. We show the resulting plots at the Figure 4. Note, that as a ground-truth we used $\theta^{*}=0.9$. We see, that given enough data, the mass of the posterior is concentrated around the correct value of the parameter, regardless of the prior misspecification.
(a) Uniform prior. $\alpha=\beta=1$.
(b) Underestimating prior. $\alpha=3,\beta=17$.
(c) Correct prior. $\alpha=9,\beta=1$.
Figure 4: Different shapes of the posterior distributions.
M.2 Prior misspecification
In this section, we demonstrate, how the metrics of epistemic uncertainty behave, given different (possibly misspecified) priors. We present the results in Figure 5. We see, that as we discussed above, these metrics depend on the values if $\alpha$ and $\beta$. Therefore, the more these values are, the less is the estimate of epistemic uncertainty. Since uniform prior contains minimal evidence, it demonstrates maximal uncertainty, and therefor less vulnerable for prior misspecification. Note, that in our experiments on deep ensembles, we considered uniform prior over parameters, therefore the effect of prior misspecification should be negligible.
(a) EPKL for different prior choices.
(b) EPBS for different prior choices.
(c) MI for different prior choices.
(d) RMI for different prior choices.
Figure 5: Epistemic uncertainty metrics, given prior misspecification and different samples sizes.
M.3 Error in posterior approximation
In this section, we consider the case of choosing wrong posterior distribution for inference. We know, that in our toy example, the correct family to consider for inference is Beta. However, we will misspecify the choice and consider Normal distribution instead. We use two approaches for inference. Namely, Laplace Approximation (approximates the posterior around the mode), and Moment Matching (matches the mean and variance of the Beta distribution with a Gaussian).
We present results in Figure 6. We considered uniform distribution as a prior, and we used training datasets of sizes 20 and 100. As expected, we see, that the given less data, errors in approximations are more severe. However, if a bigger dataset is provided, the approximation becomes more accurate.
(a) Approximations for n=20.
(b) Approximations for n=100.
Figure 6: Resulting approximations given different sizes of training datasets.